# Enforcing Paraphrase Generation via Controllable Latent Diffusion

### v1

@@ Abstract
Paraphrase generation aims to produce high-quality and diverse utterances of a given text. Though state-of-the-art generation via the diffusion model reconciles generation quality and diversity, textual diffusion suffers from a truncation issue that hinders efficiency and quality control. In this work, we propose Latent Diffusion Paraphraser (LDP), a novel paraphrase generation by modeling a controllable diffusion process given a learned latent space. LDP achieves superior generation efficiency compared to its diffusion counterparts. It facilitates only input segments to enforce paraphrase semantics, which further improves the results without external features. Experiments show that LDP achieves improved and diverse paraphrase generation compared to baselines. Further analysis shows that our method is also helpful to other similar text generations and domain adaptations. Our code and data are available at https://github.com/NIL-zhuang/ld4pg.
@@
Enforcing Paraphrase Generation via Controllable Latent Diffusion
Wei Zou [^1] Equal Contribution, Ziyuan Zhuang [^1] Equal Contribution, Shujian Huang [^2] Corresponding author, Jia Liu, Jiajun Chen National Key Laboratory for Novel Software Technology, Nanjing University Collaborative Innovation Center of Novel Software Technology and Industrialization, Nanjing {zouw, ziyuan.zhuang}@smail.nju.edu.cn {huangsj, liujia, chenjj}@nju.edu.cn
@@ 1 Introduction
Paraphrase generation aims to produce semantically equivalent sentences in varied linguistic forms. It’s versatile in generation tasks such as text summarization Zhou and Bhat (2021); Dong et al. (2017) and question answering Buck et al. (2017). Techniques such as data augmentations Kumar et al. (2019) also involve paraphrasing.
@@
Though mainstream paradigms have achieved success, they still struggle to balance generation quality and diversity. The deterministic paradigm via the encoder-decoder model Vaswani et al. (2017) focuses on high-quality generation instead of diversity. Such paradigm is further improved in diversity by enforcing explicit external features of high-level semantics shared by paraphrases, such as syntax Sun et al. (2021); Bao et al. (2019) and examplars Yang et al. (2021); Chen et al. (2019). However, external features for diversity are not always available. On the other hand, the variational paradigm promotes diversity via variational autoencoders, which model the latent distribution shared by diverse contextual representations Bao et al. (2019); Du et al. (2022). However, its sampling nature given limited Gaussian distributions hinders the generation quality of complex language patterns despite additional keyword guidance Chen et al. (2022).
@@
Recently, a novel generation paradigm via the diffusion probabilistic model (Ho et al., 2020, DPM) achieves state-of-the-art generation in both quality and diversity for images and speech. The DPM is the generation that morphs toward high-quality data through numerous rounds of continuous Markov transitions. Though DPM shares similar stochasticity with the variational generation, it does not fall short in quality. Additionally, its neat interventions enable the continuous diffusion transitions to meet the quality required by versatile data generations Nichol and Dhariwal (2021), which plays a vital role in diffusion implementation. Therefore, we consider the diffusion paradigm to fit paraphrase generation, where its controllability also meets the intuition of possible intervention in traditional paradigms.
@@
Lately, Diffusion-LM Li et al. (2022) and D3PM Austin et al. (2021) further cater to text generation via an additional discrete sampling called ‘rounding’ process. The ‘rounding’ essentially bridges the continuous diffusion representations with corresponding discrete tokens, as arbitrary diffusion intervals are truncated to embeddings of valid texts with additional decodings. However, ‘rounding’ introduces decoding overhead, thus hindering high-efficiency generation by SOTA diffusion implementations Ye et al. (2023); Karras et al. (2022). Furthermore, ‘rounding’ also introduces truncation errors when arbitrary diffusion intervals are rounded to specific token embeddings, further hindering possible intervention. Therefore, we consider circumventing the truncation issue to improve efficiency and enable generation interventions when paraphrasing via text diffusion.
@@
In this work, we propose Latent Diffusion Paraphraser (LDP), which can enforce semantics by only input segments instead of external features. LDP adopts the latent space from a given encoder-decoder framework, which offers more efficacy than raw features for diffusion as suggested by Rombach et al. (2022); Lovelace et al. (2022). The off-the-shelf encoder and decoder bridge the continuous diffusion process with corresponding discrete texts, thus LDP prevents the intermediate roundings required by diffusion on raw text, which offers generation efficiency. Furthermore, removing roundings enables state-of-the-art control for diffusion steps, where we further utilize only input segments rather than external features to enforce semantics for improvement. Experiments show that LDP achieves better and faster paraphrase generation than its diffusion counterparts on various datasets. Further analysis shows that our methods are helpful to other similar text generations and domain adaptation.
@@
  • We propose a novel paraphrase generation called LDP, which improves generation quality and diversity. LDP circumvents ‘rounding’ thus more efficient compared to its diffusion counterparts.
  • LDP can enforce paraphrase semantics with only input segments instead of external features, which further improves results.
  • Analysis shows that our method is also helpful in other similar text scenarios such as question generation and domain adaptation.
@@ 2.1 Paraphrase Generation
Paraphrase refers to the diverse utterances that keep the original semantic. Paraphrase generation is crucial in several downstream natural language processing (NLP) tasks. Though methods based on deterministic seq2seq framework have achieved success Vaswani et al. (2017); Sancheti et al. (2022); Yang et al. (2019), the nature of maximum likelihood estimation hinders the generation diversity. Some researchers promote generation diversity by enforcing explicit external features of high-level semantics shared by paraphrases such as syntactic structures or exemplar syntax Hosking et al. (2022); Yang et al. (2022), which are not easily accessible. Others turn to variational generation to fit the shared latent distribution of diverse textual representations Bowman et al. (2015); Du et al. (2022). However, variational generations sacrifice the quality for its diversity.
@@ 2.2 Latent Diffusion Models with Control
The diffusion probabilistic model (Ho et al., 2020, DPM) is a Markov chain of variational reconstruction of the original inputs $z_{0}$ given data distribution $q(z)$ from Gaussian-distributed noise. Specifically, the DPM is trained by sampling from a Markov noising process $P(z_{t+1}|z_{t})\sim\mathcal{N}(z_{t+1};\sqrt{1-\beta_{t}}z_{t},\beta_{t}\mathcal{I})$ scheduled by series of noise scales $\beta_{t}\in(0,1)$, where the $\beta_{t}$ is considered the standard deviation of the step-wise transition distribution $\sqrt{\beta_{t}}=\sigma_{t}$ at step $t\in[0,T]$. Rombach et al. further proposes the latent diffusion where the diffusion process is introduced to the latent space of a well-trained encoder-decoder framework. The latent diffusion achieves better results for video He et al. (2023), audio Liu et al. (2023), and image synthesis Lai et al. (2023) since diffusion upon encoded latent features proves more efficacy than that upon raw representations.
@@
where $z_{\theta}(\cdot)$ is a reconstruction neural net given noised $z_{t}$ and noising step $t$. Ho et al. (2020) further deducted t-steps Markov-Gaussian noising process into a one-step noising, leading to a closed form $z_{t}$ by given noising step $t$:
@@
The corresponding diffusion generation iteratively morph from pure Gaussian noise to valid data by the learned reconstructor $z_{\theta}(\cdot)$. That is, given a fully optimized $z_{\theta}(\cdot)$, the generation starts from pure Gaussian sample $z_{T}\in\mathbb{R}^{l\times d}\sim\mathcal{N}(0,I)$ with a roughly reconstructed $\hat{z}_{0}$, then iteratively noise and denoise by diminishing noise scales determined by $\beta_{t}$. Such iteration can be conducted via various diffusion sampling algorithms such as ancestral sampling Ho et al. (2020), DDIM sampler Song et al. (2020a), or ODE solversLu et al. (2022a, b).
Figure 1: Overview of LDP. (a) Model architecture. LDP consists of an encoder-decoder framework (pink) and a diffusion module (green). The encoder (E) and decoder (D) are frozen to bridge the continuous diffusion process with corresponding discrete texts. (b) Detailed architecture of denoising model $z_{\theta}(\cdot)$
The iterative diffusion generation can be intervened by plug-and-play modules with minor overheads, where the state-of-the-art modeling is ControlNet Zhang et al. (2023). ControlNet finetunes a trainable copy of the DPM $z_{\theta}^{\prime}(\cdot)$ to cater to versatile generations while freezing the learned DPM $z_{\theta}(\cdot)$ to maintain harmless adaptation. The trainable copy $z_{\theta}^{\prime}(\cdot)$ takes in additional control signals to the DPM via zero-convolution layers, i.e., convolution layers initialized by zero weight and bias. The controller inputs $\delta$ are fused with original diffusion steps by Eq 3:
$\begin{split}{z}_{t+1}=&z_{\theta}(z_{t},t)+\\ &\text{zero\_conv}(z_{\theta}^{\prime}(z_{t}+\text{zero\_conv}(\delta),t)),\end{split}$ (3)
where $\delta$ enables vague yet versatile guidance such as Canny edges and poses Zhang et al. (2023).
@@ 3 Methodology
In this section, we detail Latent Diffusion Paraphraser (LDP). LDP models the diffusion process upon the latent space of a pretrained encoder-decoder framework, where mere input segments can further enforce paraphrase semantics for improvement.
@@ 3.1 Latent Diffusion Paraphraser
The overall pipeline is shown in Fig 1(a). LDP consists of an encoder-decoder framework to provide a learned latent space and a diffusion module. The encoder and decoder with a learned latent space bridge the continuous diffusion process with corresponding discrete texts once and for all at the beginning and the end of the generation, instead of step-wise rounding. Notably, LDP is compatible with the mainstream encoder-decoder framework as long as it bijectively maps the texts with the corresponding latent representations.
@@
To make life easier, we adopt BART Lewis et al. (2019) for illustration, which is an off-the-shelf pretrained language model that encodes and decodes arbitrary text with its corresponding latent representation. Given a text sequence $x=\{x_{1},x_{2},\dots,x_{l}\}$, the BART encoder E encodes it into a $d$-dimentional latent representation $z=\text{E}(x),z\in\mathbb{R}^{l\times d}$ for diffusion process, while the decoder D yields corresponding text sequence given the latent representation $y\thickapprox\text{D}(z)=\text{D}(\text{E}(x))$. The model utilizes T5 relative positional embeddings Raffel et al. (2020). Input features for diffusion are normalized by training data features, where we adopt BART encodings for mean and standard deviation Rombach et al. (2022). Concordantly, the normalization is reversed before text decoding. The encoder and decoder are frozen during diffusion training, thus leaving the reconstruction network $z_{\theta}(\cdot)$ the only trainable parameters.
@@
Given paraphrase pair $\langle x,y\rangle$, we then train the end-to-end diffusion models for encoder latent space, where the reconstruction network is parameterized by $z_{\theta}(z_{t},c,t)$ with addition source encoding $c=E(x)$ as input. $z_{\theta}(\cdot)$ consists of $N$ layers of pre-layer norm transformer blocks. As shown in Fig 1(b), each layer consists of a self-attention encoding for $z_{t}$, followed by a cross-attention access of the encoded source $c$ and a time step interpolation via a feedforward layer. The time step $t$ is embedded as a d-dimensional vector, then its interpolation is preprocessed by AdaLN (Xu et al., 2019; Peebles and Xie, 2022, adaptive layer norm) instead of generic layer norm. AdaLN regresses the layer-wise normalization scale and shifts from the sum of time embedding and encoded input features. The network layers are activated by GeGLU Shazeer (2020) following SOTA transformer implementation Raffel et al. (2020).
@@
$z_{\theta}(\cdot)$ is optimized by reconstruction loss in Eq 1. We uniformly sample time step $t$ for arbitrary noised representation $z_{t}$ by Eq 2, where $t$ determines the noise scale $\beta_{t}$ by noise schedule Ho et al. (2020); Nichol and Dhariwal (2021). We also apply sentence-level condition dropout during training to ensure the model’s unconditional language generation, that is, to replace the source sentence representation with trainable null tokens $y_{\varnothing}$ with probability $p=0.1$.
@@
The optimized $z_{\theta}(\cdot)$ is implemented in SOTA diffusion samplers Song et al. (2020b); Lu et al. (2022a, b), which will morph pure Gaussian noise $z_{T}$ into the latent representation $\hat{z}_{0}$ of the given source text. $D$ further decodes $\hat{z}_{0}$ for text output. Unlike generation with length prediction, LDP determines the sequence length by end-of-sequence label automatically.
3.2 Semantic Enforcing by Controller
@@
Diffusion generation can introduce additional controls via ControlNet Zhang et al. (2023), which is a step-wise plug-and-play controlling framework. It has proved its efficacy by manipulating image generation given vague constraints such as canny edges or skeleton poses. By removing step-wise rounding, LDP circumvents the truncation issue against control signals, where we consider paraphrase generation can harness input segments as the vague constraints likewise.
Intuitively, the paraphrase can be enforced by mere segments of vital semantics via a controller for better generation. Therefore, we leverage input tokens above a certain length as keywords, enforcing the diffusion generation on given semantics to improve paraphrase. Our implementation is shown in Fig 2. We sample from the longest $15\%$ tokens in a given sentence as keywords and mask the remaining parts with placeholder <M> as semantic segments. The semantic segments are then encoded by the same encoder used by the original DPM for controller inputs. The controller block $z_{\theta}^{\prime}(\cdot)$ is a trainable copy of the well-trained $z_{\theta}(\cdot)$, while $z_{\theta}(\cdot)$ is frozen. The controller inputs are fused with original diffusion generation by Eq 3 as:
$\begin{split}\hat{z}_{0}=&z_{\theta}(z_{t},c,t)+\\ &\text{zero\_conv}(z_{\theta}^{\prime}(z_{t}+\text{zero\_conv}(c_{kw}),c,t)),\end{split}$
@@
QQP Twitter
[…]
Transformer 30.36 214.98 58.01 33.53 31.06 299.47 58.42 48.11
BART-FT 33.73 189.23 57.20 33.94 33.21 324.14 58.15 47.10
T5-GPVAE 33.50 200.13 64.15 25.89 27.01 129.86 58.59 52.82
[…]
DiffuSeq 24.13 397.68 86.41 49.11 9.83 1045.88 85.45 50.99
SeqDiffuSeq 24.32 404.28 70.35 48.28 11.92 903.26 71.51 52.46
[…]
Table 1: Results on paraphrase generation. Baselines are implemented from the source code. The bold indicates the top results, whereas the second best is underlined. LDP outperforms SOTA diffusion baselines. Overall, LDP with enforced semantics (LDP w/ES) achieves the best iBScore.

### v2

# Improved Paraphrase Generation via Controllable Latent Diffusion

@@ Abstract
Paraphrase generation strives to generate high-quality and diverse expressions of a given text, a domain where diffusion models excel. Though SOTA diffusion generation reconciles generation quality and diversity, textual diffusion suffers from a truncation issue that hinders efficiency and quality control. In this work, we propose Latent Diffusion Paraphraser (LDP), a novel paraphrase generation by modeling a controllable diffusion process given a learned latent space. LDP achieves superior generation efficiency compared to its diffusion counterparts. It can facilitate only input segments to ensure paraphrase semantics, improving the results without external features. Experiments show that LDP better reconciles paraphrase generation quality and diversity than baselines. Further analysis shows that our method is also helpful to other similar text generations and domain adaptations
keywords:
paraphrase generation, latent diffusion models, controlled generation
@@ 1 Introduction
Paraphrase generation aims to produce semantically equivalent sentences in varied linguistic forms. It’s versatile in generation tasks such as text summarization [1, 2] and question answering [3]. It also contributes to popular techniques such as data augmentations [4].
@@
Despite viable mainstream paradigms, traditional paraphrasing struggles to balance generation quality and diversity. The deterministic paradigm via the encoder-decoder model [5, enc-dec] focuses on high-quality generation instead of diversity. Such paradigm is further improved in diversity by enforcing explicit external features of high-level semantics shared by paraphrases, such as syntax [6] and examplars [7]. However, external features for diversity are not always available. On the other hand, the variational paradigm promotes diversity via variational autoencoders, which model the latent distribution shared by diverse contextual representations [8, 9]. However, its sampling nature given condensed Gaussian distributions hinders the generation quality of complex language patterns despite additional keyword guidance [10].
@@
Images and speech recently witnessed a novel generation paradigm via the diffusion probabilistic model [11, DPM] achieving SOTA in both quality and diversity. The DPM is the generation that morphs toward high-quality data through numerous rounds of continuous Markov transitions. Though DPM shares similar stochasticity with the variational generation, it does not fall short in quality. Additionally, its neat interventions enable the continuous diffusion transitions to meet the quality required by versatile data generations [12], which is prominent in diffusion implementation. However, DPM works in a continuous space, which differs from the discrete text token.
@@
Lately, Diffusion-LM [13] and D3PM [14] cater to text generation via an additional discrete sampling called ‘rounding’ process. The ‘rounding’ essentially bridges the continuous diffusion representations with corresponding discrete tokens, as arbitrary diffusion intervals are truncated to embeddings of valid texts with additional decodings. However, ‘rounding’ introduces decoding overhead, thus hindering high-efficiency generation by SOTA diffusion implementations [15, 16]. Furthermore, ‘rounding’ also introduces truncation errors when arbitrary diffusion intervals are rounded to specific token embeddings, further hindering possible intervention. Therefore, we consider circumventing the truncation issue to improve efficiency and enable generation interventions when paraphrasing via text diffusion.
@@
In this work, we propose the Latent Diffusion Paraphraser (LDP), which not only enjoys the benefit of DPM in balancing quality and diversity but also utilizes flexible control in diffusion modeling for paraphrasing. LDP adopts the latent space from a given enc-dec framework, which offers more efficacy than raw features for diffusion as suggested by work [17, 18]. The off-the-shelf encoder and decoder bridge the continuous diffusion process with corresponding discrete texts. Thus, LDP prevents the intermediate roundings required by diffusion on raw text, which offers generation efficiency. Furthermore, removing roundings enables SOTA diffusion controls, where we can even utilize only input segments rather than external features to ensure semantics for improvement. Experiments show that LDP achieves better and faster paraphrase generation than its diffusion counterparts on various datasets. Further analysis shows that our methods are viable to other similar text generations and domain adaptation.
@@
  • We propose a novel paraphrase generation called LDP, which better reconciles generation quality and diversity elegantly and efficiently.
  • LDP utilizes the flexible diffusion control on semantics for improvement even with only input segments rather than external features.
  • Analysis shows that our method is also helpful in similar text scenarios such as question generation and domain adaptation.
@@ 2.1 Paraphrase Generation
Paraphrase refers to the diverse utterances that keep the original semantic. Paraphrase generation is crucial in several downstream natural language processing (NLP) tasks. Though methods based on deterministic seq2seq framework achieve success [5, 19, 20], the nature of maximum likelihood estimation hinders the generation diversity. Some researchers promote generation diversity by enforcing explicit external features of high-level semantics shared by paraphrases such as syntactic structures or exemplar syntax [21, 22], which are not easily accessible. Others turn to variational generation to fit the shared latent distribution of diverse textual representations [23, 9]. However, variational generations sacrifice the quality for its diversity.
@@ 2.2 Latent Diffusion Models with Control
The diffusion probabilistic model [11, DPM] is a Markov chain of variational reconstruction of the original inputs $z_{0}$ given data distribution $q(z)$ from Gaussian-distributed noise. Specifically, the DPM is trained by sampling from a Markov noising process $P(z_{t+1}|z_{t})\sim\mathcal{N}(z_{t+1};\sqrt{1-\beta_{t}}z_{t},\beta_{t}\mathcal{I})$ scheduled by series of noise scales $\beta_{t}\in(0,1)$, where the $\beta_{t}$ is considered the standard deviation of the step-wise transition distribution $\sqrt{\beta_{t}}=\sigma_{t}$ at step $t\in[0,T]$. Work [17] further proposes the latent diffusion where the diffusion process is introduced to the latent space of a well-trained enc-dec framework. The latent diffusion achieves better results for video [24], audio [25], and image synthesis [26] since diffusion upon encoded latent features proves more efficacy than that upon raw representations.
@@
where $z_{\theta}(\cdot)$ is a reconstruction neural net given noised $z_{t}$ and noising step $t$. Work [11] further deducted t-steps Markov-Gaussian noising process into a one-step noising, leading to a closed form $z_{t}$ by given noising step $t$:
@@
The corresponding diffusion generation iteratively morphs from pure Gaussian noise to valid data by the learned reconstructor $z_{\theta}(\cdot)$. That is, given a fully optimized $z_{\theta}(\cdot)$, the generation starts from pure Gaussian sample $z_{T}\in\mathbb{R}^{l\times d}\sim\mathcal{N}(0,I)$ with a roughly reconstructed $\hat{z}_{0}$, then iteratively noise and denoise by diminishing noise scales determined by $\beta_{t}$. Such iteration can be conducted via various diffusion sampling algorithms such as ancestral sampling [11], DDIM sampler [27], or ODE solvers[28, 29].
Figure 1: Overview of LDP. (a) Model architecture. LDP consists of an encoder-decoder (enc-dec) framework (pink) and a diffusion module (green). The encoder (E) and decoder (D) are frozen to bridge the continuous diffusion process with corresponding discrete texts. Where the source $x$ is encoded as $c$ during the diffusion on latent variable $z$. (b) Detailed architecture of denoising model $z_{\theta}(\cdot)$.
Plug-and-play modules can flexibly intervene in iterative diffusion generation with minor overheads, where the SOTA model is ControlNet [30]. ControlNet is an adaptor that finetunes a trainable copy of the DPM $z_{\theta}^{\prime}(\cdot)$ to cater to versatile generations while freezing the learned DPM $z_{\theta}(\cdot)$ to maintain harmless adaptation. The trainable copy $z_{\theta}^{\prime}(\cdot)$ takes in additional control signals to the DPM via zero-convolution layers, i.e., convolution layers initialized by zero weight and bias. The controller inputs $\delta$ are fused with original diffusion steps by Eq 3:
$\begin{split}{z}_{t+1}=&z_{\theta}(z_{t},t)+\text{zero\_conv}(z_{\theta}^{\prime}(z_{t}+\text{zero\_conv}(\delta),t)),\end{split}$ (3)
where $\delta$ enables vague yet versatile guidance such as Canny edges and poses [30].
@@ 3 Methodology
In this section, we detail Latent Diffusion Paraphraser (LDP). LDP models the diffusion process upon the latent space of a pretrained enc-dec framework, where mere input segments can further guide paraphrase semantics for improvement.
@@ 3.1 Latent Diffusion Paraphraser
The overall pipeline is shown in Fig 1(a). LDP consists of an enc-dec framework to provide a learned latent space and a diffusion module. The encoder and decoder with a learned latent space bridge the continuous diffusion process with corresponding discrete texts once and for all at the beginning and the end of the generation, instead of step-wise rounding. Notably, LDP is compatible with the mainstream enc-dec framework as long as it bijectively maps the texts with the corresponding latent representations.
@@
To make life easier, we adopt BART [31] for illustration, which is an off-the-shelf pretrained language model that encodes and decodes arbitrary text with its corresponding latent representation. Given a text sequence $x=\{x_{1},x_{2},\dots,x_{l}\}$, the BART encoder E encodes it into a $d$-dimentional latent representation $z=\text{E}(x),z\in\mathbb{R}^{l\times d}$ for diffusion process, while the decoder D yields corresponding text sequence given the latent representation $y\thickapprox\text{D}(z)=\text{D}(\text{E}(x))$. The model utilizes T5 relative positional embeddings [32]. Input features for diffusion are normalized by training data features, where we adopt BART encodings for mean and standard deviation [17]. Concordantly, the normalization is reversed before text decoding. The encoder and decoder are frozen during diffusion training, thus leaving the reconstruction network $z_{\theta}(\cdot)$ the only trainable parameters.
@@
Given paraphrase pair $\langle x,y\rangle$, we then train the end-to-end diffusion models for encoder latent space, where the reconstruction network is parameterized by $z_{\theta}(z_{t},c,t)$ with addition source encoding $c=E(x)$ as input. $z_{\theta}(\cdot)$ consists of $N$ layers of pre-layer norm transformer blocks. As shown in Fig 1(b), each layer consists of a self-attention encoding for $z_{t}$, followed by a cross-attention access of the encoded source $c$ and a time step interpolation via a feedforward layer. The time step $t$ is embedded as a d-dimensional vector, then its interpolation is preprocessed by AdaLN [33, 34, adaptive layer norm] instead of the generic layer norm. AdaLN regresses the layer-wise normalization scale and shifts from the sum of time embedding and encoded input features. The network layers are activated by GeGLU [35] following SOTA transformer implementation [32].
@@
$z_{\theta}(\cdot)$ is optimized by reconstruction loss in Eq 1. We uniformly sample time step $t$ for arbitrary noised representation $z_{t}$ by Eq 2, where $t$ determines the noise scale $\beta_{t}$ by noise schedule [11, 12]. We also apply sentence-level condition dropout during training to ensure the model’s unconditional language generation, that is, to replace the source sentence representation with trainable null tokens $y_{\varnothing}$ with probability $p=0.1$.
@@
The optimized $z_{\theta}(\cdot)$ is implemented in SOTA diffusion samplers [36, 28, 29], which will morph pure Gaussian noise $z_{T}$ into the latent representation $\hat{z}_{0}$ of the given source text. $D$ further decodes $\hat{z}_{0}$ for text output. Unlike generation with length prediction, LDP determines the sequence length by end-of-sequence label automatically.
3.2 Semantics Guidance by Controller
@@
Diffusion generation can introduce additional controls via ControlNet [30], which is a step-wise plug-and-play adaptor framework. It has proved its efficacy by manipulating image generation given vague constraints such as canny edges or skeleton poses. By removing step-wise rounding, LDP circumvents the truncation issue against control signals, where we consider paraphrase generation can harness input segments as the vague constraints likewise.
Intuitively, the paraphrase can be guided by such a framework based on vital semantic fragments for better generation. We leverage input tokens above a certain length as keywords, guiding the diffusion generation on given semantics to improve paraphrase. Our implementation is shown in Fig 2. We sample from the longest $15\%$ tokens in a given sentence as keywords and mask the remaining parts with placeholder <M> as semantic segments. The semantic segments are then encoded by the same encoder used by the original DPM for controller inputs. The controller block $z_{\theta}^{\prime}(\cdot)$ is a trainable copy of the well-trained $z_{\theta}(\cdot)$, while $z_{\theta}(\cdot)$ is frozen. The controller inputs are fused with original diffusion generation by Eq 3 as:
$\begin{split}\hat{z}_{0}=&z_{\theta}(z_{t},c,t)+\text{zero\_conv}(z_{\theta}^{\prime}(z_{t}+\text{zero\_conv}(c_{kw}),c,t)),\end{split}$
@@
QQP WIKI Twitter
BLEU PPL div-4
iBScore
(BScore-srcBLEU)
BLEU PPL div-4
iBScore
(BScore-srcBLEU)
BLEU PPL div-4
iBScore
(BScore-srcBLEU)
Transformer 30.36 214.98 58.01
33.53
(86.92-55.39)
50.64 202.60 68.32
13.15
(89.91-76.76)
31.06 299.47 58.42
48.11
(81.90-33.79)
BART-FT 33.73 189.23 57.20
33.94
(54.32-88.26)
62.79 112.51 61.67
10.87
(93.89-83.02)
33.21 324.14 58.15
47.10
(81.74-34.64)
T5-GPVAE 33.50 200.13 64.15
25.89
(87.84-61.95)
61.05 109.07 63.98
10.89
(89.91-76.76)
27.01 129.86 58.59
52.82
(79.83-27.01)
BART-CVAE 32.21 194.81 55.40
47.02
(87.56-40.54)
61.14 116.88 64.16
9.68
(93.81-84.13)
32.24 325.71 59.77
44.37
(81.48-37.11)
DiffuSeq 31.49 397.68 86.41
49.11
(85.89-36.78)
25.89 440.49 80.67
48.35
(84.47-36.12)
9.83 1045.88 85.45
54.61
(75.06-20.46)
SeqDiffuSeq 26.67 404.28 70.35
48.28
(84.56-36.28)
28.13 309.88 73.13
13.15
(89.91-76.76)
11.92 903.26 71.51
52.46
(76.62-24.16)
Llama2-7b 16.93 218.02 66.22
50.10
(85.95-35.85)
23.26 166.70 62.09
60.73
(86.74-26.01)
4.28 262.46 60.65
57.47
(73.74-16.27)
Llama3.1-8b 13.90 210.21 67.23
57.97
(85.37-27.40)
16.56 187.29 60.67
63.31
(85.35-23.03)
3.42 240.44 57.87
61.48
(75.25-13.77)
LDP(ours) 36.56 267.53 73.22
50.57
(87.09-36.52)
60.38 120.13 81.2
58.15
(85.15-27.00)
18.75 466.46 93.32
59.72
(76.14-16.42)
w/ SG 37.48 246.76 73.63
51.44
(88.82-37.40)
61.99 146.54 72.22
64.91
(91.34-26.76)
19.72 419.63 93.26
60.36
(82.05-21.69)
Table 1: Results on paraphrase generation. Baselines are implemented from the source code. The bold indicates the top results, whereas the second bests are underlined. We group the SOTA baselines into two groups, the traditional and diffusion baselines. LDP outperforms SOTA diffusion baselines of traditional baselines. Notably, open-source LLMs have a model size of 30-40$\times$ larger than LDP. LDP with semantics guidance (LDP w/SG) evan achieves the iBScore equivalent to open-source LLMs, with significantly higher BLEU scores.